On-Time Airplane Arrivals. The greater Cincinnati airport led major U.S. airports in on-time arrivals in the last quarter of 2005 with an 84.3% on-time rate. choose 5 arrivals at random and find the probablitiy that at least 1 was not on time.
Answer by stanbon(75887)

On-Time Airplane Arrivals. The greater Cincinnati airport led major U.S. airports in on-time arrivals in the last quarter of 2005 with an 84.3% on-time rate. choose 5 arrivals at random and find the probability that at least 1 was not on time.
----------------------
Binomial Problem with n = 5 and p(on-time) = 0.843 ; p(not on-time) = 0.157
-------------------
P(1<= not on time <=5) = 1 - P(x = 0 not on time)
-----
= 1 - 0.843^5
----
= 1 - 0.5743
----
= 0.4247
==============================
Cheers,
Stan H.
==============